### v1

# Just-in-time and distributed task representations in language models

Abstract
Many of language models’ impressive capabilities originate from their in-context learning: based on instructions or examples, they can infer and perform new tasks without weight updates. In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context. We focus on “transferrable” task representations—vector representations that can restore task context in another instance of the model, even without the full prompt. We show that these representations evolve in non-monotonic and sporadic ways, and are distinct from a more inert representation of high-level task categories that persists throughout the context. Specifically, models often condense multiple evidence into these transferrable task representations, which align well with the performance improvement based on more examples in the context. However, this accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context. Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask, and models rely on more temporally-distributed representations to support longer and composite tasks. This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process underlying language models’ ability to adapt to new evidence and learn new tasks on the fly.
1 Introduction
Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; Brown et al.,, 2020): the emergent ability to learn tasks from few-shot examples in context. This discovery has led to a variety of works exploring the behavioral features of ICL (e.g. Sclar et al.,, 2024; Min et al.,, 2022). Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (Agarwal et al.,, 2024; Anil et al.,, 2024). The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities. Work on interpreting the mechanisms responsible for ICL has led to discoveries such as induction heads (e.g. Olsson et al.,, 2022) and how ICL implicitly refines a model of in-context evidence (e.g. Akyürek et al.,, 2022; Von Oswald et al.,, 2023).
Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (Todd et al.,, 2024; Hendel et al.,, 2023). Importantly, these task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt. This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream—such that task context is reinstantiated and the model can perform the task without any explicit instructions or demonstrations. These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024).
The discovery of transferrable task representations raises several intriguing questions about their dynamics and generality. How and when are these internal task representations formed throughout the context? How do these dynamics depend upon task complexity? A simple, intuitive hypothesis is that these representations develop gradually, e.g. at a rate that depends on task complexity. The representations might accrue evidence from each example and refine monotonically into a more stable, robust task representation. This view aligns with the behavioral findings that models perform better with more examples in-context (Agarwal et al.,, 2024; Anil et al.,, 2024).
We set out to understand how the dynamics of ICL (within and across examples) are exhibited representationally within the model, by investigating task representations. Our findings suggest a more complicated picture of the computational process that support language models’ adaptation to a new task:
  • For simple tasks, we find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
  • These fleeting but transferrable task representations often condense evidence from multiple examples but tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
  • Models do not appear to condense task knowledge into local task representations in more complicated tasks that require state tracking or chaining multiple subtasks together.
  • Finally, models can form distinct representations to support generating the same responses when solving tasks independently vs. as part of a broader context.
Overall, these results give us a window into language models’ changing state when inferring and solving new tasks in context, but paint a complex and nuanced picture of the dynamics of this state. There are different types of task representations—identifiable vs. transferrable—that evolve over the context in distinct ways. The representations of tasks also depend on the task complexity and the surrounding context structure in which a task is embedded. These results may have implications for both the science of understanding models, and practical applications of mechanistic interpretability for analysis and safety.
Figure 1: Understanding how task representations develop over context. A. A schematic of extracting transferrable task representations and restoring task contexts in zero-shot settings. The highlighted tokens indicate the source and target for extracting and injecting task representations. B. Transferrable task representations restore task accuracy on zero-shot prompts. Results are aggregated over all models for simple tasks (see Appendix A). Error bars indicate the 95% CI over tasks. C. An overview of the development of transferrable task representations over context, i.e. the recontextualized zero-shot accuracy for task representations extracted from different tokens throughout the context.
2 Methods
Tasks
For our analyses, we built upon tasks from prior work on transferrable task representations (Hendel et al.,, 2023; Todd et al.,, 2024). The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word, translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words). In addition to these simple, single-token generation tasks from the previous literature, we also test a range of new tasks to explore model behavior in longer generation settings. These include: repeating a simple task three times (e.g. antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose the first and last words from a word list), and reversing or shifting an entire word list. Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each word in a word list. (See Appendix A for the full set of tasks.)
There are 512 query-answer pairs for each task (except for two smaller datasets: country-capital contains 197 samples, and product-company contains 494 samples). These query-answer pairs are formatted into few-shot prompts with alternating "Q:" and "A:" turns, as shown in Figure 1A.
Models
We focus our analyses on the open-weight Gemma V3 pre-trained models, including the 4B, 12B, and 27B-sized models (Team et al.,, 2025).
Extracting in-context task representations
We primarily investigate task vectors discovered in (Hendel et al.,, 2023) as a window to study language models’ in-context task representations. We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query as shown in Figure 1A. Task vectors are the layer residual activations extracted from the last token before answer generation in the few-shot prompts (in the example in Figure 1A, this corresponds to the highlighted colon). Hendel et al., (2023) showed that task vectors can reinstate task performance even without any additional context and with a different query—when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any few-shot examples in-context.
We replicate and extend the procedure outlined in Hendel et al., (2023). For each model and task, we first search for the layer that best captures the task representation, using 50 queries from the dataset as the development set and in an 8-shot setting. As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations. We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model. The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task, and subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset. Consistent with prior results, we generally find that task vectors extracted and injected at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
Evaluating task accuracy
We compare the average accuracy for the sampled responses across a few settings: standard zero-shot, recontextualized zero-shot setting (with task vector intervention), and few-shot. For simplicity, responses for all tasks are graded by exact string matches against the ground-truth answer. This underestimates the model performance in some tasks (e.g. for antonym and translation tasks), but we use the same grading scheme across all settings and compare relative performances. For longer-generation and mixed-generation tasks, we evaluate each of the multiple outputs separately by exact match (e.g., in antonym x 3, we compare each of the three output words with the correct answer), and report the mean accuracy across all outputs in the longer response.
Examining the dynamics of task representations
Once we determine the best layer for each task using the last colon token, we evaluate how well the colon token representations condense the k-shot information in a prompt. We do this by extracting task vectors at the colon token for a k-shot prompt, and then evaluating the recontextualized zero-shot accuracy when the task vectors are patched into a zero-shot prompt. We repeat this analysis for k in 1, 2, 4, 8, 16, 32. We also experimented with allowing the best layer to vary depending on k, but found very similar results.
We also studied whether transferrable task representations form in tokens other than the final colon token. In particular, we extracted layer residual token activations for other tokens in the context, including the "Q", the ":" following "Q", the "A", and the new-line token before "A". We patched these token activations onto the corresponding token in the zero-shot prompt at the same layer. For each of the non-final tokens, we repeated the search for the layer that best captures task representations. All token representations were evaluated on the extent to which they restored task accuracy on zero-shot prompts.
3 Results
How do language models infer new tasks in-context? We leverage the ability for in-context task representations to restore task contexts to understand how models accrue evidence and refine task representations. We find that language models indeed form stronger task representations that aggregates in-context evidence. However, this evidence accrual process is surprisingly non-gradual and happens in a sporadic way for most tasks (Figure 1C). In particular, effective and transferrable task representations form only at certain tokens in the context, and tend to capture the minimal “task scope.” For some tasks, models also do not appear to condense task information into local task representations. Below, we discuss these findings in more detail.
Figure 2: Sporadic & inconsistent evidence accrual in language models. A. Task vectors extracted from the last colon token in each example capture evidence accrual on most tasks (11 out of 14). However, on three “hard-to-transfer” tasks, task vectors do not capture this evidence accumulation, even though the models (behaviorally) do learn from more examples. The solid bars indicate recontextualized zero-shot accuracy (via task vectors), and light bars in the background indicate 8-shot accuracy (without task vectors). The dotted lines indicate the ratio of the recontextualized zero-shot accuracy against 8-shot accuracy. B. Most other format tokens in the context do not robustly form transferrable task representations that support recontextualization on zero-shot, but task identity is reliably decodable in their residual activations. Here, we report the task identity decoding accuracy at the mode best layer at which transferrable task representations form in the second ":" token. See the main text and Appendix C for more details.
3.1 Local task representations can accrue evidence
Task representations can reflect evidence accrual.
Consistent with the behavioral performance gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations in models also reflect increased task certainty with increased context. As shown in Figure 2A, task vectors extracted from more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples. This suggests that language models condense information from multiple examples and form better task representations, even when the task representations extracted are fairly local (i.e., at the last token in the prompt).
… but not for all types of tasks.
However, we did not observe evidence accrual in 3 out of 14 of the simple tasks (Figure 2A, hard-to-transfer tasks). Specifically, for choose_middle_of_5, count_color_in_3, and count_fruit_in_3 tasks, we found that local task representations (extracted from the last token) were not able to take advantage of more examples for task transfer to zero-shot settings, even though models improved substantially at solving the task when given more examples in the prompt. One interpretation is that some tasks require more state-tracking (such as counting), and that these may necessitate additional inference processes that the models do not condense into local task representations. Alternatively, these inference processes cannot be effectively re-activated by the injection of the extracted task representations.
How evidence accrual leads task representations to converge.
For the eleven other simple tasks where we successfully observed evidence accrual, we sought to understand how the task representations themselves changed over more examples (Figure 3A). As we increased the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot samples. This suggests that, as models gain evidence, in-context task representations tend to denoise or converge to more stable representations. The magnitude (L2-norm) of the task vectors also tends to decrease over time. However, there were noticeable differences between tasks and models. Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2). However, for certain tasks, the magnitude of the task representations first increases then decreases given more examples. We note that for this analysis, we look at the task representations extracted from the mode best layer across different tasks. This is to control for magnitude difference of the residual activations across layers and make a fair comparison. Although the best layer for transferrable task representations does sometimes differ across tasks, the best layer across tasks tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
Figure 3: Analyses of extracted task representations. A. The extracted task vectors (at the last colon token) decrease in both variance and magnitude with more examples, exhibiting a general tendency to condense evidence and converge onto stable task representations. The solid line shows the average across tasks, the transparent lines show the individual tasks. B. Extracted task vectors from the 27B model form distinct clusters. The numbers label the centroid for each task (see legend and results for other models in Figure S4). Task vectors are similar but distinguishable when a task was evaluated independently vs. embedded within a larger task structure, e.g., representations for antonym (0), antonym x 3 (14), and where antonym appears as a first task in a mixed-generation task chain (24&27) are close but distinct.
3.2 Task representations exhibit temporal locality
The analyses above confirm that, intuitively, in-context task representations can successfully benefit from increasing evidence in context (along the “temporal” dimension), and converge to better representations that restore task contexts more robustly. To understand the full temporal profile of this accrual process, we repeated the task vector recontextualization experiment on other tokens in the prompt, which revealed that the transferrable task representations do not strengthen monotonically.
Transferrable task representations are not found in most tokens.
We extracted task vectors from the format tokens in the prompt, including "Q", the ":" following "Q", "A", and the new-line token before the "A". We tested whether these extracted representations can also restore the corresponding task contexts when patched onto the forward pass in the zero-shot prompt (as before, we patched at the same layer, but onto the corresponding format token instead of the last colon). As shown in Figure 2B, transferrable task representations generally do not robustly form in the residual activations across layers in these tokens, even though these tokens are also shared across examples, tasks, and contexts. This is true across the number of examples provided in the prompt, leading to the developmental trajectory of in-context task representations shown in Figure 1C.
We observed nearly zero recontextualized zero-shot accuracy for all these tokens in most tasks, except some restoration success in product-company, color_v_animal_3, choose_first_of_5, and the longer-generation tasks discussed below (see Figure S3). In general, it seems that an effective, transferrable task representation in language models only forms sparingly; in few-shot settings, this often means a just-in-time task representation at the token just before answer generation for each query.
… but a robust task identity signal persists throughout the context.
Intriguingly, however, task identity is almost perfectly decodable in the representations extracted from all the different format tokens, even though the formats are shared across the tasks and contexts. To study this, we trained simple, single-layer linear decoders to predict the task category (among the eleven simple tasks that we observed evidence accrual) from the residual activations of each format token. We report the test accuracy on the 25% held-out token activations extracted following eight examples in Figure 2B and Appendix C, but found high task identity decoding accuracy across these tokens following different numbers of examples. This decodability success may be partly due to vocabulary differences for the tasks, but the accuracy is close to ceiling, and not all tasks are distinguishable in this way. We note that task identity is not perfectly decodable in the token activations in early layers, but generally reaches high accuracy in mid and late layers. This suggests that the model is generally task-sensitive, but instantiates transferrable task representations only at particular timepoints in the context.
3.3 Task representations exhibit scope locality
Figure 4: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over longer generation, especially for tasks that can be decomposed into semantically-independent subtasks. This suggests a tendency for models to only activate transferrable representations for small task scopes. A. Left: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks. Right: recontextualized accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context. An output unit usually corresponds to a single word and is occasionally a short phrase (e.g. the capital of a country). B. Visualization as in A, but for mixed-generation tasks.
We have seen evidence that transferrable task representations tend to be temporally local. That leads to the question of whether they have a lasting effect over generation – that is, are the restored task contexts in the zero-shot forward pass also fleeting in nature? Prior work has mostly focused on simple, single-token output tasks. To study the semantic scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token. Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
Transferrable task representations tend to support a semantically-independent task scope.
In these experiments, we find further evidence of the locality of in-context task representations. Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 4, left; also see Figure S1). Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 4, right), suggesting that the restored task contexts “fade” over longer generation.
Models seem to be consistently decomposing tasks and forming local task representations that capture a minimal “task scope”. For example, in choose_first_middle_last_of_5, the reinstantiated task context only supported generating the first word, suggesting that further token generations rely on additional representations instantiated by a separate mechanism while processing the previously-generated tokens. This effect is very pronounced in the mixed-generation tasks, as all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain and defer representing later subtasks. Interestingly, the extracted task representations that support generating responses for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 3B). This can potentially reflect the inert sensitivity to the broader mixed-task context, as well as conveying information relevant for later subtasks to be activated, even though these signals are not sufficient to directly restore execution of later subtasks.
The restored task contexts do support longer generation to some extent in some tasks, including repeating a simple task multiple times and the list-level tasks (i.e. reversing or shifting a word list). Overall, these results suggest that language models may be automatically segmenting semantically-independent task scopes when possible, such that task representations for longer or composite tasks are offloaded onto multiple tokens.
4 Related work
Since the discovery that large language models exhibit emergent in-context learning (Brown et al.,, 2020), there has been substantial interest in investigating this capability and its mechanistic basis.
From a behavioral perspective, many subsequent works have explored how ICL could develop from implicit meta-learning of data properties (Xie et al.,, 2022; Chan et al.,, 2022), and how this may relate to the broader set of language model capabilities (Chen et al.,, 2024; Lampinen et al.,, 2024). Some of this work has focused on the surprising fragility of ICL to subtle prompt changes (e.g. Sclar et al.,, 2024); conversely, others have highlighted how ICL may be overly robust, allowing “learning” common tasks even if the labels are randomized (Min et al.,, 2022). One particularly relevant focus of behavioral work on ICL has been on the dynamics of in-context learning; for example, how adding many example shots can improve performance on difficult tasks, or even those discouraged in post-training (Agarwal et al.,, 2024; Anil et al.,, 2024).
From a mechanistic perspective, Olsson et al., (2022) showed how induction heads could support in-context learning, and other work has studied how they might develop over training (Edelman et al.,, 2024; Singh et al.,, 2025). More recently, attention has turned to cases where models may create internal task representations, that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance. Hendel et al., (2023) demonstrated an instance of such task representations: representations at intermediate layers of the model at a key token that can be injected to mimic the effect of a few-shot prompt. Concurrent work from Todd et al., (2024) identified “function vectors,” which aggregate the effects of multiple attention heads that convey task information. Subsequent work has generalized and extended these findings, for example exploring how function vectors can emerge from instructions as well as examples (Davidson et al.,, 2025) and how task vectors capture task representations in multimodal models (Huang et al.,, 2024; Luo et al.,, 2024). Other works have explored how these representations and structures emerge over training (Yang et al.,, 2025; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024). Our work builds on these findings and use transferrable task contexts as a window to study the dynamics of in-context task representations in language models.
5 Discussion
We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks. Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task context in zero-shot settings. Our results show that in many tasks, models refine task representations over more evidence such that the representations more successfully restore task contexts. However, these transferrable task representations only sporadically activate, and seem to best support a minimal task scope (e.g., a first irreducible subtask). The dynamics of these effective, transferrable task representations strongly contrast with a general sensitivity to high-level task differences that persists throughout the context. In addition, we find many cases where models do not appear to condense global task representations into a local, token-level representation, such as tasks involving more state tracking and tasks chaining different types of subtasks together.
In general, our results complicate the intuitive picture that the computational process underlying language model’s in-context evidence accrual and task inference is smooth and gradual. Models do not appear to refine task representations at a per-token basis, not even in a step-wise manner—even on a per-example basis, the task state is not sustained, but often fades and reactivates across different tokens. That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information at multiple layers may suggest a general inductive bias to compress knowledge to local representations when possible. This may also relate to the success many works have observed on extracting transferrable task representations in broad settings, such as following instructions or images (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024), or capturing information for multiple possible task outcomes (Xiong et al.,, 2024).
However, one trend that arose from these investigations is that language models also do not seem able to condense task information into local representations in all cases. Rather, they exhibit a tendency to form sharp local contexts for small task units, and offload broader task contexts such as mixed or multiple tasks across time. As discussed earlier, even for simple tasks where some intermediate-state tracking may be required, successful inference may need to rely on cross-token and cross-layer computation (e.g., as shown in Ameisen et al.,, 2025). In these cases, the effective restoration of the computation process may also require intervening multiple components during the forward pass. It is also possible that by overriding token activations with task vectors in intermediate layers, the models may have lost any task state information formed in earlier layers, and other methods that restore task states through additive injection rather than activation patching may be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024). Understanding how more complicated in-context computation develops based on evidence in these tasks and whether vector-form task representations can re-activate these inferences are interesting open questions.
An intriguing direction for future work is to study whether there are mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations. One possibility may be that the residual stream is more stable and easier to learn from during training. This may encourage the model to rely more on the residual stream to condense contextual task representations rather than relying on the more expensive attention operations. These learning dynamics may drive models to conform with an implicit normative consideration to not instantiate task contexts until needed and instantiate just the right scope to avoid capacity waste. Some of these features may even be exclusive to models pre-trained on natural languages (e.g. Yang et al.,, 2025).
Limitation
We note a few important limitations of our work. First, we primarily rely on task vectors as the method to extract transferrable task representations. This means that our conclusions and speculations are bounded by the effectiveness of this method. As we discussed earlier, representations for some task contexts may be more distributed, either across tokens and/or across model layers. It would be important to confirm if similar dynamics are observed in less-constrained methods such as a multi-layer recontextualization (Li et al.,, 2024). Second, we mostly explored relatively simple tasks, including when we investigated longer-generation tasks. It’s possible that many of the dynamics we observe here would not generalize to settings with naturalistic languages, especially when the tasks are not so cleanly decomposable and a single, semantically-independent task unit is hard to define.
Conclusion
We investigated how the dynamics of in-context learning are reflected in the development of language models’ internal task representations. Our results suggest that language models do not smoothly refine a global task state in-context. While general task sensitivity persists throughout context, models appear to construct effective task representations in a “just-in-time” fashion to solve a simple task scope immediately ahead. The fleeting, minimally-scoped nature of these in-context task representations provides new insight into the models’ state of inferring and performing tasks based on new evidence.
Acknowledgments and Disclosure of Funding
We thank Yasaman Bahri, Katherine Hermann, Mike Mozer, Sjoerd van Steenkiste, and Murray Shanahan for helpful comments and support on this work.
Appendix A Tasks
Table 1: Simple/shorter-answer tasks. See Todd et al., (2024) for more details for the first nine tasks.
Task Name Example
antonym Q: true A: false
country-capital Q: Germany A: Berlin
english-french Q: queens A: reines
product-company Q: Windows XP A: Microsoft
color_v_animal_3 Q: blue, dolphin, swan A: blue
fruit_v_animal_3 Q: lime, parrot, buffalo A: lime
choose_first_of_5 Q: envelope, pasta, cake, toucan, create A: envelope
choose_middle_of_5 Q: candy, charismatic, laptop, realize, eel A: laptop
choose_last_of_5 Q: affable, believe, carefree, zoom, moray A: moray
word_length Q: negotiate A: 9
count_color_in_3 Q: snake, gold, indigo A: two
count_fruit_in_3 Q: lime, newt, bunny A: one
position_of_color_in_3 Q: monkey, oryx, white A: third
position_of_fruit_in_3 Q: pear, coyote, capybara A: first
Table 2: Longer-generation tasks.
Task Name Example
antonym x 3 Q: fall, everybody, intact A: rise, nobody, broken
product-company x 3 Q: iWork, Windows NT 3.5, OS X Yosemite A: Apple, Microsoft, Apple
all_but_color_in_3 Q: cat, black, pelican A: cat, pelican
all_but_fruit_in_3 Q: grape, butterfly, llama A: butterfly, llama
choose_first_last_of_3 Q: white, house, wallet A: white, wallet
choose_first_middle_last_of_5 Q: dolphin, beyond, curtain, pillow, intuitive A: dolphin, curtain, intuitive
reverse_all_of_3 Q: donut, sad, who A: who, sad, donut
reverse_all_of_5 Q: she, honest, out, test, frog A: frog, test, out, honest, she
shift_all_of_3 Q: piano, cougar, jackfruit A: cougar, jackfruit, piano
shift_all_of_5 Q: agreeable, flamingo, short, around, jovial A: flamingo, short, around, jovial, agreeable
Table 3: Mixed-generation tasks.
Task Name Example
antonym + product-company Q: opponent, iDisk A: ally, Apple
english-french + antonym Q: liberal, continue A: libéral, stop
product-company + english-french Q: Alfa Romeo MiTo, mask A: Fiat, masque
antonym + country-capital + english-french Q: upper, Greece, artists A: lower, Athens, artistes
english-french + product-company + antonym Q: system, Lancia Flavia, unlucky A: système, Fiat, lucky
country-capital + antonym + product-company Q: Gambia, heavy, Game & Watch A: Banjul, light, Nintendo
Appendix B Additional figures
Figure S1: Recontextualization accuracy for all tasks. Task vectors extracted from 8-shot prompts are used to reinstantiate task contexts in zero-shot settings. The dotted line indicate the ratio between recontextualized zero-shot accuracy and 8-shot accuracy.
Figure S2: Developmental trajectory of task representations over shots. Task representations are the token activations of the colon token prior to answer generation. We visualize task vectors sourced from the mode best layer across tasks at which task contexts are best restored in a zero-shot setting. Representations are first averaged across samples with the same number of examples for each task.
Figure S3: Recontextualized zero-shot accuracy from different format tokens in the prompt. The colors indicate different models (see Figure S1).
Figure S4: Task representations across all models and tasks.
Appendix C Additional results
Table 4: Decoding accuracy for task identity at different tokens. Results shown for token activations extracted at the 8th example.
[…]
model layer
GEMMA_V3_4B_PT 2 0.933 0.848 0.925 0.899 0.950
5 0.979 0.980 0.816 0.925 0.994
8 0.952 0.986 0.908 0.959 0.996
11 0.950 0.986 0.968 0.984 0.998
14 0.960 0.990 0.988 0.984 0.996
17 0.902 0.994 0.990 0.996 0.996
20 0.918 0.992 0.970 0.982 0.992
23 0.834 0.988 0.992 0.954 0.996
26 0.842 0.865 0.990 0.916 0.992
29 0.811 0.973 0.992 0.922 0.986
32 0.905 0.948 0.992 0.958 0.966
GEMMA_V3_12B_PT 2 0.945 0.896 0.858 0.808 0.963
5 0.886 0.965 0.872 0.959 0.905
8 0.975 0.894 0.866 0.889 0.979
11 0.958 0.940 0.980 0.998 0.986
14 0.980 0.960 0.980 0.988 0.998
17 0.979 0.986 0.956 0.946 0.998
20 0.977 0.990 0.968 0.942 0.998
23 0.963 0.998 0.994 0.933 0.984
26 0.907 1.000 0.998 0.848 0.992
29 0.889 0.994 0.982 0.859 1.000
32 0.924 0.962 0.984 0.742 1.000
35 0.845 0.956 0.952 0.869 0.994
38 0.773 0.942 0.980 0.808 0.986
41 0.756 0.869 0.920 0.772 0.990
44 0.817 0.928 0.984 0.953 0.990
47 0.956 0.971 0.970 0.928 0.968
GEMMA_V3_27B_PT 5 0.929 0.876 0.829 0.879 0.905
11 0.940 0.994 0.865 0.946 0.972
17 0.979 0.990 0.900 0.996 0.996
23 0.973 1.000 0.945 0.960 0.996
29 0.950 0.998 0.994 0.937 0.994
35 0.864 0.988 0.924 0.906 0.994
41 0.936 0.964 0.960 0.827 0.990
47 0.890 0.996 0.912 0.858 0.990
53 0.887 0.980 0.912 0.787 0.968
59 0.863 0.984 0.994 0.977 0.958

### v2

@@
 In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context.
-We focus on “transferrable” task representations—vector representations that can restore task context in another instance of the model, even without the full prompt.
+We focus on “transferrable” task representations—vector representations that can restore task contexts in another instance of the model, even without the full prompt.
 We show that these representations evolve in non-monotonic and sporadic ways, and are distinct from a more inert representation of high-level task categories that persists throughout the context.
-Specifically, models often condense multiple evidence into these transferrable task representations, which align well with the performance improvement based on more examples in the context.
-However, this accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context.
-Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask, and models rely on more temporally-distributed representations to support longer and composite tasks.
-This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process underlying language models’ ability to adapt to new evidence and learn new tasks on the fly.
+Specifically, when more examples are provided in the context, transferrable task representations successfully condense evidence.
+This allows better transfer of task contexts and aligns well with the performance improvement.
+However, this evidence accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context.
+Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask.
+For longer and composite tasks, models rely on more temporally-distributed representations.
+This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process that language models use to perform new tasks on the fly.
 1 Introduction
@@
 Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (Agarwal et al.,, 2024; Anil et al.,, 2024).
-The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities.
-Work on interpreting the mechanisms responsible for ICL has led to discoveries such as induction heads (e.g.
+The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities, leading to discoveries such as induction heads (e.g.
 Olsson et al.,, 2022) and how ICL implicitly refines a model of in-context evidence (e.g.
@@
 Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (Todd et al.,, 2024; Hendel et al.,, 2023).
-Importantly, these task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt.
-This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream—such that task context is reinstantiated and the model can perform the task without any explicit instructions or demonstrations.
+These task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt.
+This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream.
+Such intervention reinstates the task context and allows the model to perform the task without any explicit instructions or demonstrations.
 These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024).
@@
 at a rate that depends on task complexity.
-The representations might accrue evidence from each example and refine monotonically into a more stable, robust task representation.
+The representations might accrue evidence across tokens and refine monotonically into a more stable and robust task representation.
 This view aligns with the behavioral findings that models perform better with more examples in-context (Agarwal et al.,, 2024; Anil et al.,, 2024).
-We set out to understand how the dynamics of ICL (within and across examples) are exhibited representationally within the model, by investigating task representations.
+We set out to investigate these transferrable task representations to understand how the dynamics of ICL are exhibited within them.
 Our findings suggest a more complicated picture of the computational process that support language models’ adaptation to a new task:
-  • For simple tasks, we find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
-  • These fleeting but transferrable task representations often condense evidence from multiple examples but tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
+  • We find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
+  • The fleeting but transferrable task representations condense evidence from multiple examples.
+However, they also tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
   • Models do not appear to condense task knowledge into local task representations in more complicated tasks that require state tracking or chaining multiple subtasks together.
@@
 C.
-An overview of the development of transferrable task representations over context, i.e.
-the recontextualized zero-shot accuracy for task representations extracted from different tokens throughout the context.
+An overview of the development of different task representations over context.
+Solid bars: recontextualized zero-shot accuracy for task vectors extracted from different tokens.
+Transparent bars: task identity decoding accuracy from different token representations.
 2 Methods
@@
 For our analyses, we built upon tasks from prior work on transferrable task representations (Hendel et al.,, 2023; Todd et al.,, 2024).
-The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word, translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words).
+The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word or translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words).
 In addition to these simple, single-token generation tasks from the previous literature, we also test a range of new tasks to explore model behavior in longer generation settings.
 These include: repeating a simple task three times (e.g.
-antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose the first and last words from a word list), and reversing or shifting an entire word list.
-Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each word in a word list.
-(See Appendix A for the full set of tasks.)
+antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose both the first and the last word in the list), and reversing or shifting an entire word list.
+Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each input item.
+See Appendix A for the full set of tasks.
 There are 512 query-answer pairs for each task (except for two smaller datasets: country-capital contains 197 samples, and product-company contains 494 samples).
-These query-answer pairs are formatted into few-shot prompts with alternating "Q:" and "A:" turns, as shown in Figure 1A.
+These query-answer pairs are formatted into few-shot prompts with alternating “Q:” and “A:” turns, as shown in Figure 1A.
 Models
-We focus our analyses on the open-weight Gemma V3 pre-trained models, including the 4B, 12B, and 27B-sized models (Team et al.,, 2025).
-Extracting in-context task representations
-We primarily investigate task vectors discovered in (Hendel et al.,, 2023) as a window to study language models’ in-context task representations.
-We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query as shown in Figure 1A.
-Task vectors are the layer residual activations extracted from the last token before answer generation in the few-shot prompts (in the example in Figure 1A, this corresponds to the highlighted colon).
-Hendel et al., (2023) showed that task vectors can reinstate task performance even without any additional context and with a different query—when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any few-shot examples in-context.
+In the main paper, we present results on the open-weight pre-trained 4B, 12B, and 27B Gemma V3 models (Team et al.,, 2025).
+In Appendix C, we also show that the main findings replicate on the 4B, 8B, and 14B Qwen3 models (Yang et al., 2025a, ).
+Extracting transferrable task representations
+We primarily investigate task vectors discovered in Hendel et al., (2023) as a window to study language models’ transferrable task representations.
+We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query, as shown in Figure 1A.
+Task vectors are the layer residual activations extracted from the last token before answer generation (in the example in Figure 1A, this corresponds to the highlighted colon token).
+Hendel et al., (2023) showed that task vectors can reinstate task performance on a different query even without any prior context.
+Specifically, when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any prior few-shot examples.
 We replicate and extend the procedure outlined in Hendel et al., (2023).
@@
 As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations.
-We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model.
-The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task, and subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset.
+We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model (covering the global attention layers).
+The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task.
+This best layer is subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset.
 Consistent with prior results, we generally find that task vectors extracted and injected at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
-Evaluating task accuracy
-We compare the average accuracy for the sampled responses across a few settings: standard zero-shot, recontextualized zero-shot setting (with task vector intervention), and few-shot.
+Evaluating task transfer
+We compare the average accuracy of the sampled responses across three settings: standard zero-shot, recontextualized zero-shot (with task vector intervention), and few-shot (with examples in context).
 For simplicity, responses for all tasks are graded by exact string matches against the ground-truth answer.
@@
 for antonym and translation tasks), but we use the same grading scheme across all settings and compare relative performances.
-For longer-generation and mixed-generation tasks, we evaluate each of the multiple outputs separately by exact match (e.g., in antonym x 3, we compare each of the three output words with the correct answer), and report the mean accuracy across all outputs in the longer response.
-Examining the dynamics of task representations
-Once we determine the best layer for each task using the last colon token, we evaluate how well the colon token representations condense the k-shot information in a prompt.
-We do this by extracting task vectors at the colon token for a k-shot prompt, and then evaluating the recontextualized zero-shot accuracy when the task vectors are patched into a zero-shot prompt.
-We repeat this analysis for k in 1, 2, 4, 8, 16, 32.
-We also experimented with allowing the best layer to vary depending on k, but found very similar results.
-We also studied whether transferrable task representations form in tokens other than the final colon token.
-In particular, we extracted layer residual token activations for other tokens in the context, including the "Q", the ":" following "Q", the "A", and the new-line token before "A".
+For longer-generation and mixed-generation tasks, we evaluate each of the multiple outputs separately by exact match (e.g., in antonym x 3, we compare each of the three output words with the correct answer), and report the average accuracy across all output units.
+Examining the dynamics of transferrable task representations
+Once we determine the best layer for each task using the last colon token, we evaluate how well the colon token representations condense information from multiple examples in a prompt.
+We do this by extracting task vectors at the colon token in different k-shot prompts and patching onto zero-shot prompts, then comparing the recontextualized zero-shot accuracy.
+We repeat this analysis for k in 0, 1, 2, 4, 8, 16, 32.
+In an earlier experiment on a subset of the tasks, we also experimented with allowing the best layer to vary depending on k, but found very similar results overall.
+We show these layer search results across different k’s in Figure 2A, but otherwise focus on results from reusing the best layer from the layer search on 8-shot prompts for other k’s.
+We then studied whether transferrable task representations form in tokens other than the colon token.
+We extracted layer residual token activations for other format tokens in the context, including the “Q”, the “:” following “Q”, the “A”, and the new-line token before “A”.
 We patched these token activations onto the corresponding token in the zero-shot prompt at the same layer.
-For each of the non-final tokens, we repeated the search for the layer that best captures task representations.
-All token representations were evaluated on the extent to which they restored task accuracy on zero-shot prompts.
+For each of the non-colon tokens, we repeat the search for the layer that best captures task representations.
+All token representations are evaluated on the extent to which they restore task accuracy on zero-shot prompts.
+Decoding task identity
+We also studied whether token representations throughout the context contain robust task identity information.
+We trained simple linear decoders to predict the task category from the layer residual activations of different tokens.
+At each layer and token combination, we used 100 token instances for each task to train and test a task identity decoder.
+All decoders were trained for 20 epochs, with a batch size of 256 and a learning rate of 0.01.
+We report the decoding accuracy across the 25% held-out test representations.
 3 Results
 How do language models infer new tasks in-context?
-We leverage the ability for in-context task representations to restore task contexts to understand how models accrue evidence and refine task representations.
-We find that language models indeed form stronger task representations that aggregates in-context evidence.
-However, this evidence accrual process is surprisingly non-gradual and happens in a sporadic way for most tasks (Figure 1C).
-In particular, effective and transferrable task representations form only at certain tokens in the context, and tend to capture the minimal “task scope.” For some tasks, models also do not appear to condense task information into local task representations.
-Below, we discuss these findings in more detail.
+We leverage the transferability of task contexts to understand how models accrue evidence and refine task representations.
+We find that language models indeed form stronger task representations that aggregate in-context evidence.
+However, this evidence accrual process is non-gradual and happens in a sporadic way.
+In particular, models only form effective task representations that transfer task contexts at certain tokens (Figure 1C, also see Figure 2A).
+This is in strong contrast to a persistent sensitivity of task identity across tokens in the context (Figure 2B).
+We also find that the surprisingly local transferrable task representations tend to only capture a minimal “task scope.” Below, we discuss these findings in more detail.
+Figure 2: Transferrable task representations activate sporadically at key tokens, but task identity representations persist throughout the context.
+A.
+Recontextualization accuracy when each token representation is used to restore task contexts in zero-shot settings.
+B.
+Task identity decoding accuracy (among 14 tasks) for token representations at different layers and positions.
+This figure plots aligned sequences across different samples and tasks; since exact positions differ depending on the sample, the indices shown in the labels are approximate.
+See results for other models in Figure S4.
 Figure 3: Sporadic & inconsistent evidence accrual in language models.
 A.
-Task vectors extracted from the last colon token in each example capture evidence accrual on most tasks (11 out of 14).
-However, on three “hard-to-transfer” tasks, task vectors do not capture this evidence accumulation, even though the models (behaviorally) do learn from more examples.
-The solid bars indicate recontextualized zero-shot accuracy (via task vectors), and light bars in the background indicate 8-shot accuracy (without task vectors).
-The dotted lines indicate the ratio of the recontextualized zero-shot accuracy against 8-shot accuracy.
+Task vectors extracted from the last colon token in each example capture evidence accrual on most tasks (12 out of 14).
+However, on two “hard-to-transfer” tasks, task vectors do not capture this evidence accumulation, even though the models (behaviorally) do learn from more examples.
+The solid bars indicate recontextualized zero-shot accuracy (via task vectors), and light bars in the background indicate few-shot accuracy (without task vectors).
+The dotted lines indicate the ratio of the recontextualized zero-shot accuracy against few-shot accuracy.
 B.
 Most other format tokens in the context do not robustly form transferrable task representations that support recontextualization on zero-shot, but task identity is reliably decodable in their residual activations.
-Here, we report the task identity decoding accuracy at the mode best layer at which transferrable task representations form in the second ":" token.
-See the main text and Appendix C for more details.
+Here, we report the task identity decoding accuracy at the mode best layer at which transferrable task representations form in the second “:” token.
+See the main text for more details.
 3.1 Local task representations can accrue evidence
-Task representations can reflect evidence accrual.
-Consistent with the behavioral performance gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations in models also reflect increased task certainty with increased context.
-As shown in Figure 2A, task vectors extracted from more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples.
-This suggests that language models condense information from multiple examples and form better task representations, even when the task representations extracted are fairly local (i.e., at the last token in the prompt).
+Transferrable task representations reflect evidence accrual.
+Consistent with the behavioral gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations also reflect increased task certainty with increased context (Figure 3A).
+Task vectors extracted following more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples.
+This suggests that language models condense information from multiple examples and form better task representations, even though the task representations extracted are fairly local (i.e., from a single token in the few-shot prompts).
 … but not for all types of tasks.
-However, we did not observe evidence accrual in 3 out of 14 of the simple tasks (Figure 2A, hard-to-transfer tasks).
-Specifically, for choose_middle_of_5, count_color_in_3, and count_fruit_in_3 tasks, we found that local task representations (extracted from the last token) were not able to take advantage of more examples for task transfer to zero-shot settings, even though models improved substantially at solving the task when given more examples in the prompt.
-One interpretation is that some tasks require more state-tracking (such as counting), and that these may necessitate additional inference processes that the models do not condense into local task representations.
+However, we did not observe strong evidence accrual in 2 out of the 14 simple tasks (Figure 3A, hard-to-transfer tasks).
+These two tasks are count_color_in_3 and count_fruit_in_3.
+For all three model sizes, task vectors extracted from 32-shot prompts recovered below 50% of the few-shot accuracy with 32 examples in the context.
+In other words, local task representations for these tasks were not able to take advantage of more examples for task transfer to zero-shot settings, even though models improved substantially at solving the task when given more examples in the prompt.
+Interestingly, the same counting tasks were hard-to-transfer for Qwen3 models as well (see Figure S7A).
+One possibility is that these tasks require more state-tracking, which may necessitate additional inference processes that the models do not condense into local task representations.
 Alternatively, these inference processes cannot be effectively re-activated by the injection of the extracted task representations.
-How evidence accrual leads task representations to converge.
-For the eleven other simple tasks where we successfully observed evidence accrual, we sought to understand how the task representations themselves changed over more examples (Figure 3A).
-As we increased the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot samples.
-This suggests that, as models gain evidence, in-context task representations tend to denoise or converge to more stable representations.
+How evidence accrual leads transferrable task representations to converge.
+As models appear to successfully accrue evidence in these highly local representations, we sought to understand how the task representations themselves changed over more examples (Figure 4A).
+For this analysis, we look at the task representations extracted from the mode best layer across different tasks.
+This is to control for magnitude differences of the residual activations across layers and make a fair comparison.
+Although the best layer for transferrable task representations sometimes differ across tasks, the best layers tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
+In Gemma V3 models, as we increase the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot prompts.
+This can signal that in-context task representations tend to denoise or converge to more stable representations as models gain evidence.
 The magnitude (L2-norm) of the task vectors also tends to decrease over time.
-However, there were noticeable differences between tasks and models.
+However, for both the variance and magnitude, there were considerable differences between tasks and model sizes.
 Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2).
-However, for certain tasks, the magnitude of the task representations first increases then decreases given more examples.
-We note that for this analysis, we look at the task representations extracted from the mode best layer across different tasks.
-This is to control for magnitude difference of the residual activations across layers and make a fair comparison.
-Although the best layer for transferrable task representations does sometimes differ across tasks, the best layer across tasks tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
-Figure 3: Analyses of extracted task representations.
+For certain tasks, the magnitude of the task representations first increases then decreases given more examples.
+We also note that both the variance and magnitude of task vectors tend to increase per more examples in some Qwen3 models (Figure S8), suggesting that different models may develop different strategies for refining task representations.
+Figure 4: Analyses of extracted task representations in Gemma V3 models.
 A.
-The extracted task vectors (at the last colon token) decrease in both variance and magnitude with more examples, exhibiting a general tendency to condense evidence and converge onto stable task representations.
-The solid line shows the average across tasks, the transparent lines show the individual tasks.
+The extracted task vectors (at the last colon token) tend to decrease in both variance and magnitude with more examples, exhibiting a general tendency to condense evidence and converge onto stable task representations.
+The solid line shows the average across tasks.
+The transparent lines show the individual tasks.
 B.
@@
 The numbers label the centroid for each task (see legend and results for other models in Figure S5).
-Task vectors are similar but distinguishable when a task was evaluated independently vs.
-embedded within a larger task structure, e.g., representations for antonym (0), antonym x 3 (14), and where antonym appears as a first task in a mixed-generation task chain (24&27) are close but distinct.
+Task vectors are similar but distinguishable when a task is evaluated independently vs.
+embedded within a larger task structure.
+For example, representations for antonym (0), antonym x 3 (14), and where antonym appears as a first task in a mixed-generation task chain (24&27) are close but distinct.
 3.2 Task representations exhibit temporal locality
-The analyses above confirm that, intuitively, in-context task representations can successfully benefit from increasing evidence in context (along the “temporal” dimension), and converge to better representations that restore task contexts more robustly.
+The analyses above confirm that transferrable task representations benefit from increasing in-context evidence and converge to better representations.
 To understand the full temporal profile of this accrual process, we repeated the task vector recontextualization experiment on other tokens in the prompt, which revealed that the transferrable task representations do not strengthen monotonically.
 Transferrable task representations are not found in most tokens.
-We extracted task vectors from the format tokens in the prompt, including "Q", the ":" following "Q", "A", and the new-line token before the "A".
-We tested whether these extracted representations can also restore the corresponding task contexts when patched onto the forward pass in the zero-shot prompt (as before, we patched at the same layer, but onto the corresponding format token instead of the last colon).
-As shown in Figure 2B, transferrable task representations generally do not robustly form in the residual activations across layers in these tokens, even though these tokens are also shared across examples, tasks, and contexts.
-This is true across the number of examples provided in the prompt, leading to the developmental trajectory of in-context task representations shown in Figure 1C.
+We tested whether extracted representations from other format tokens in the prompt can also restore the corresponding task contexts.
+These include “Q”, the “:” following “Q”, “A”, and the new-line token before “A”, which are all shared across examples, tasks, and contexts.
+As before, we patched the activations at the same layer, but onto the corresponding format token instead of the last colon token in zero-shot prompts.
+As shown in Figure 3B, transferrable task representations generally do not form in the residual activations across layers in these tokens.
+This is true across the number of examples provided in the prompt, leading to the developmental trajectory of transferrable task representations shown in Figure 1C.
 We observed nearly zero recontextualized zero-shot accuracy for all these tokens in most tasks, except some restoration success in product-company, color_v_animal_3, choose_first_of_5, and the longer-generation tasks discussed below (see Figure S3).
-In general, it seems that an effective, transferrable task representation in language models only forms sparingly; in few-shot settings, this often means a just-in-time task representation at the token just before answer generation for each query.
+In Qwen3 models, there also appear to be some restoration successes with representations extracted from the “Q” token (see Figure S7B and Figure S11).
+These partial successes are likely driven by the fact that the first answer token sometimes match the first input token across a subset of the tasks.
+In general, it seems that an effective, transferrable task representation in language models only forms sparingly; in few-shot settings, this often means a just-in-time task representation at the token before answer generation.
 … but a robust task identity signal persists throughout the context.
-Intriguingly, however, task identity is almost perfectly decodable in the representations extracted from all the different format tokens, even though the formats are shared across the tasks and contexts.
-To study this, we trained simple, single-layer linear decoders to predict the task category (among the eleven simple tasks that we observed evidence accrual) from the residual activations of each format token.
-We report the test accuracy on the 25% held-out token activations extracted following eight examples in Figure 2B and Appendix C, but found high task identity decoding accuracy across these tokens following different numbers of examples.
-This decodability success may be partly due to vocabulary differences for the tasks, but the accuracy is close to ceiling, and not all tasks are distinguishable in this way.
-We note that task identity is not perfectly decodable in the token activations in early layers, but generally reaches high accuracy in mid and late layers.
+Intriguingly, however, task identity is almost perfectly decodable in the representations extracted from all the different format tokens, even though the formats are shared across all tasks.
+We report the decoder accuracy from the layer with the best transferrable task representations in Figure 3B, but found high task identity decoding accuracy across most tokens throughout the context (Figure 2B).
+The decodability success may be partly due to vocabulary differences between some tasks, but we show that decoding accuracy remains high even for a restricted subset of tasks with shared vocabulary (Figure S4C).
+Both task identity and task transferability do not occur until at least one full example is presented, but accurate task identity representations form much earlier than transferrable task representations (Figure 1C).
 This suggests that the model is generally task-sensitive, but instantiates transferrable task representations only at particular timepoints in the context.
 3.3 Task representations exhibit scope locality
-Figure 4: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over longer generation, especially for tasks that can be decomposed into semantically-independent subtasks.
+We have seen evidence that transferrable task representations tend to be temporally local.
+That leads to the question of whether they have a lasting effect over generation.
+That is, are the restored task contexts in the zero-shot forward pass also fleeting in nature?
+Prior work has mostly focused on simple, single-token output tasks.
+To study the scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token.
+Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
+Figure 5: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over generation, especially for tasks that can be decomposed into semantically-independent subtasks.
 This suggests a tendency for models to only activate transferrable representations for small task scopes.
 A.
-Left: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks.
-Right: recontextualized accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context.
+Bar plot: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks.
+Line plots: recontextualization accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context.
 An output unit usually corresponds to a single word and is occasionally a short phrase (e.g.
@@
 Visualization as in A, but for mixed-generation tasks.
-We have seen evidence that transferrable task representations tend to be temporally local.
-That leads to the question of whether they have a lasting effect over generation – that is, are the restored task contexts in the zero-shot forward pass also fleeting in nature?
-Prior work has mostly focused on simple, single-token output tasks.
-To study the semantic scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token.
-Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
 Transferrable task representations tend to support a semantically-independent task scope.
-In these experiments, we find further evidence of the locality of in-context task representations.
-Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 4, left; also see Figure S1).
-Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 4, right), suggesting that the restored task contexts “fade” over longer generation.
+In these experiments, we find further evidence of the locality of transferrable task representations.
+Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 5, bar plots; also see Figure S1).
+Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 5, line plots; see similar results on Qwen3 models in Figure S13), suggesting that the restored task contexts “fade” over longer generation.
 Models seem to be consistently decomposing tasks and forming local task representations that capture a minimal “task scope”.
 For example, in choose_first_middle_last_of_5, the reinstantiated task context only supported generating the first word, suggesting that further token generations rely on additional representations instantiated by a separate mechanism while processing the previously-generated tokens.
-This effect is very pronounced in the mixed-generation tasks, as all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain and defer representing later subtasks.
-Interestingly, the extracted task representations that support generating responses for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 3B).
-This can potentially reflect the inert sensitivity to the broader mixed-task context, as well as conveying information relevant for later subtasks to be activated, even though these signals are not sufficient to directly restore execution of later subtasks.
-The restored task contexts do support longer generation to some extent in some tasks, including repeating a simple task multiple times and the list-level tasks (i.e.
-reversing or shifting a word list).
+This effect is very pronounced in the mixed-generation tasks.
+In these cases, all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain, and defer representing later subtasks.
+Interestingly, the extracted task representations for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 4B), even though they support generating the same responses.
+This can potentially reflect an inert sensitivity to the broader mixed-task context based on vocabulary differences.
+Additionally, these differences may also convey information relevant for later subtasks to be activated, even though they are not sufficient to directly restore execution of later subtasks.
+The restored task contexts do support longer generation to some extent in many tasks, including repeating a simple task multiple times and some list-level tasks (i.e., reversing or shifting a word list).
 Overall, these results suggest that language models may be automatically segmenting semantically-independent task scopes when possible, such that task representations for longer or composite tasks are offloaded onto multiple tokens.
@@
 From a mechanistic perspective, Olsson et al., (2022) showed how induction heads could support in-context learning, and other work has studied how they might develop over training (Edelman et al.,, 2024; Singh et al.,, 2025).
-More recently, attention has turned to cases where models may create internal task representations, that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance.
+More recently, attention has turned to cases where models may create internal task representations.
+These are representations that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance.
 Hendel et al., (2023) demonstrated an instance of such task representations: representations at intermediate layers of the model at a key token that can be injected to mimic the effect of a few-shot prompt.
@@
 Subsequent work has generalized and extended these findings, for example exploring how function vectors can emerge from instructions as well as examples (Davidson et al.,, 2025) and how task vectors capture task representations in multimodal models (Huang et al.,, 2024; Luo et al.,, 2024).
-Other works have explored how these representations and structures emerge over training (Yang et al.,, 2025; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024).
+Other works have explored how these representations and structures emerge over training (Yang et al., 2025b, ; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024).
 Our work builds on these findings and use transferrable task contexts as a window to study the dynamics of in-context task representations in language models.
@@
 We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks.
-Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task context in zero-shot settings.
+Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task contexts in zero-shot settings.
 Our results show that in many tasks, models refine task representations over more evidence such that the representations more successfully restore task contexts.
@@
 In addition, we find many cases where models do not appear to condense global task representations into a local, token-level representation, such as tasks involving more state tracking and tasks chaining different types of subtasks together.
-In general, our results complicate the intuitive picture that the computational process underlying language model’s in-context evidence accrual and task inference is smooth and gradual.
+In general, our results complicate the intuitive picture that the computational process underlying language models’ in-context learning and task inference is smooth and gradual.
 Models do not appear to refine task representations at a per-token basis, not even in a step-wise manner—even on a per-example basis, the task state is not sustained, but often fades and reactivates across different tokens.
-That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information at multiple layers may suggest a general inductive bias to compress knowledge to local representations when possible.
+That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information may suggest a general inductive bias to compress knowledge to local representations when possible.
 This may also relate to the success many works have observed on extracting transferrable task representations in broad settings, such as following instructions or images (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024), or capturing information for multiple possible task outcomes (Xiong et al.,, 2024).
-However, one trend that arose from these investigations is that language models also do not seem able to condense task information into local representations in all cases.
+However, one trend that arose from these investigations is that language models do not seem able to condense task information into local representations in all cases.
 Rather, they exhibit a tendency to form sharp local contexts for small task units, and offload broader task contexts such as mixed or multiple tasks across time.
@@
 In these cases, the effective restoration of the computation process may also require intervening multiple components during the forward pass.
-It is also possible that by overriding token activations with task vectors in intermediate layers, the models may have lost any task state information formed in earlier layers, and other methods that restore task states through additive injection rather than activation patching may be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024).
-Understanding how more complicated in-context computation develops based on evidence in these tasks and whether vector-form task representations can re-activate these inferences are interesting open questions.
-An intriguing direction for future work is to study whether there are mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations.
+It is also possible that by overriding token activations at intermediate layers with task vectors, the models have lost any task state information formed in earlier layers.
+Other methods that restore task states through additive injection rather than activation patching may therefore be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024).
+We also note that different models may differ in the extent to which they form local or distributed effective representations across the context or layers.
+An intriguing direction for future work is to study the mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations.
 One possibility may be that the residual stream is more stable and easier to learn from during training.
@@
 Some of these features may even be exclusive to models pre-trained on natural languages (e.g.
-Yang et al.,, 2025).
+Yang et al., 2025b, ).
 Limitation
@@
 country-capital + antonym + product-company Q: Gambia, heavy, Game & Watch A: Banjul, light, Nintendo
-Appendix B Additional figures
+Appendix B Additional Gemma3 results
 Figure S1: Recontextualization accuracy for all tasks.
 Task vectors extracted from 8-shot prompts are used to reinstantiate task contexts in zero-shot settings.
-The dotted line indicate the ratio between recontextualized zero-shot accuracy and 8-shot accuracy.
+The dotted line indicates the ratio between recontextualized zero-shot accuracy and 8-shot accuracy.
 Figure S2: Developmental trajectory of task representations over shots.
-Task representations are the token activations of the colon token prior to answer generation.
+Task vectors are the token activations of the colon token prior to answer generation.
 We visualize task vectors sourced from the mode best layer across tasks at which task contexts are best restored in a zero-shot setting.
-Representations are first averaged across samples with the same number of examples for each task.
-Figure S3: Recontextualized zero-shot accuracy from different format tokens in the prompt. The colors indicate different models (see Figure S1).
-Figure S4: Task representations across all models and tasks.
-Appendix C Additional results
-Table 4: Decoding accuracy for task identity at different tokens. Results shown for token activations extracted at the 8th example.
-[…]
-model layer
-GEMMA_V3_4B_PT 2 0.933 0.848 0.925 0.899 0.950
-5 0.979 0.980 0.816 0.925 0.994
-8 0.952 0.986 0.908 0.959 0.996
-11 0.950 0.986 0.968 0.984 0.998
-14 0.960 0.990 0.988 0.984 0.996
-17 0.902 0.994 0.990 0.996 0.996
-20 0.918 0.992 0.970 0.982 0.992
-23 0.834 0.988 0.992 0.954 0.996
-26 0.842 0.865 0.990 0.916 0.992
-29 0.811 0.973 0.992 0.922 0.986
-32 0.905 0.948 0.992 0.958 0.966
-GEMMA_V3_12B_PT 2 0.945 0.896 0.858 0.808 0.963
-5 0.886 0.965 0.872 0.959 0.905
-8 0.975 0.894 0.866 0.889 0.979
-11 0.958 0.940 0.980 0.998 0.986
-14 0.980 0.960 0.980 0.988 0.998
-17 0.979 0.986 0.956 0.946 0.998
-20 0.977 0.990 0.968 0.942 0.998
-23 0.963 0.998 0.994 0.933 0.984
-26 0.907 1.000 0.998 0.848 0.992
-29 0.889 0.994 0.982 0.859 1.000
-32 0.924 0.962 0.984 0.742 1.000
-35 0.845 0.956 0.952 0.869 0.994
-38 0.773 0.942 0.980 0.808 0.986
-41 0.756 0.869 0.920 0.772 0.990
-44 0.817 0.928 0.984 0.953 0.990
-47 0.956 0.971 0.970 0.928 0.968
-GEMMA_V3_27B_PT 5 0.929 0.876 0.829 0.879 0.905
-11 0.940 0.994 0.865 0.946 0.972
-17 0.979 0.990 0.900 0.996 0.996
-23 0.973 1.000 0.945 0.960 0.996
-29 0.950 0.998 0.994 0.937 0.994
-35 0.864 0.988 0.924 0.906 0.994
-41 0.936 0.964 0.960 0.827 0.990
-47 0.890 0.996 0.912 0.858 0.990
-53 0.887 0.980 0.912 0.787 0.968
-59 0.863 0.984 0.994 0.977 0.958
+Representations for each task are first averaged across samples with the same number of examples in the prompt.
+Figure S3: Recontextualized zero-shot accuracy from different format tokens in the prompt in different tasks.
+The colors indicate different model sizes: blue=GEMMA_V3_4B_PT, yellow=GEMMA_V3_12B_PT, red=GEMMA_V3_27B_PT.
+Figure S4: The development of different task representations over layers and context.
+A.
+Recontextualized accuracy when each token representation is used to restore task context in zero-shot settings.
+B.
+Task identity decoding accuracy across all 14 simple/short-answer tasks in Table 1.
+C.
+Task identity decoding accuracy across 9 list operation tasks with shared vocabulary.
+This subset of tasks includes: choose_first_of_5, choose_middle_of_5, choose_last_of_5, choose_first_last_of_3, choose_first_middle_last_of_5, reverse_all_of_3, reverse_all_of_5, shift_all_of_3, shift_all_of_5.
+Figure S5: T-SNE projection of all task vectors across models and tasks.
+Appendix C Results on Qwen3 models
+Figure S6: Recontextualization accuracy from Qwen3 models for all tasks. Visualization as in Figure S1.
+Figure S7: In Qwen3 models, transferrable task representations also condense evidence and activate sporadically at specific tokens.
+Visualization as in Figures 3A, B and Figure 1C.
+The hard-to-transfer tasks in Qwen3 models include english-french, choose_first_of_5, count_color_v_animal_3, and count_fruit_v_animal_3.
+These are the tasks where task vectors extracted from 32-shot prompts failed to recover more than 50% of the 32-shot accuracy on 0-shot prompts, across all model sizes.
+Figure S8: Variance and magnitude changes across task vectors extracted following different number of examples. Visualization as in Figure 4A.
+Figure S9: T-SNE task representations across all models and tasks.
+Figure S10: Developmental trajectory of task representations over shots. Visualization as in Figure S2.
+Figure S11: Recontextualized zero-shot accuracy from different format tokens in the prompt. The colors indicate different model sizes: blue=QWEN3_4B, yellow=QWEN3_8B, red=QWEN3_14B.
+Figure S12: Task identity decoding accroacy across all 14 simple tasks in Table 1 (A) and 9 list operation tasks with shared vocabulary (B). See Figure S4 caption for a complete list of tasks.
+Figure S13: In Qwen3, restored task contexts in longer- and mixed-generation tasks also decay over more output units during generation. Visualization as in Figure 5.

### v3

@@
 In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context.
-We focus on “transferrable” task representations—vector representations that can restore task contexts in another instance of the model, even without the full prompt.
-We show that these representations evolve in non-monotonic and sporadic ways, and are distinct from a more inert representation of high-level task categories that persists throughout the context.
-Specifically, when more examples are provided in the context, transferrable task representations successfully condense evidence.
-This allows better transfer of task contexts and aligns well with the performance improvement.
-However, this evidence accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context.
-Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask.
-For longer and composite tasks, models rely on more temporally-distributed representations.
-This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process that language models use to perform new tasks on the fly.
+We study two different task representations: those that are “transferrable”—vector representations that can transfer task contexts to another model instance, even without the full prompt—and simpler representations of high-level task categories.
+We show that transferrable task representations evolve in non-monotonic and sporadic ways, while task identity representations persist throughout the context.
+Specifically, transferrable task representations exhibit a two-fold locality.
+They successfully condense evidence when more examples are provided in the context.
+But this evidence accrual process exhibits strong temporal locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context.
+In some cases, transferrable task representations also show semantic locality, capturing a small task “scope” such as an independent subtask.
+Language models thus represent new tasks on the fly through both an inert, sustained sensitivity to the task and an active, just-in-time representation to support inference.
 1 Introduction
-Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; Brown et al.,, 2020): the emergent ability to learn tasks from few-shot examples in context.
+Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; brown2020language): the emergent ability to learn tasks from few-shot examples in context.
 This discovery has led to a variety of works exploring the behavioral features of ICL (e.g.
-Sclar et al.,, 2024; Min et al.,, 2022).
-Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (Agarwal et al.,, 2024; Anil et al.,, 2024).
+sclar2024quantifying; min2022rethinking).
+Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (agarwal2024many; anil2024many).
 The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities, leading to discoveries such as induction heads (e.g.
-Olsson et al.,, 2022) and how ICL implicitly refines a model of in-context evidence (e.g.
-Akyürek et al.,, 2022; Von Oswald et al.,, 2023).
-Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (Todd et al.,, 2024; Hendel et al.,, 2023).
+olsson2022context), the core circuits responsible for learning from in-context examples (cho2024revisiting; bakalova2025contextualize), and how ICL implicitly refines a model of in-context evidence (e.g.
+akyurek2022learning; von2023transformers).
+Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (todd2023function; hendel2023context).
 These task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt.
@@
 Such intervention reinstates the task context and allows the model to perform the task without any explicit instructions or demonstrations.
-These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024).
-The discovery of transferrable task representations raises several intriguing questions about their dynamics and generality.
-How and when are these internal task representations formed throughout the context?
-How do these dynamics depend upon task complexity?
-A simple, intuitive hypothesis is that these representations develop gradually, e.g.
-at a rate that depends on task complexity.
-The representations might accrue evidence across tokens and refine monotonically into a more stable and robust task representation.
-This view aligns with the behavioral findings that models perform better with more examples in-context (Agarwal et al.,, 2024; Anil et al.,, 2024).
-We set out to investigate these transferrable task representations to understand how the dynamics of ICL are exhibited within them.
-Our findings suggest a more complicated picture of the computational process that support language models’ adaptation to a new task:
-  • We find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
-  • The fleeting but transferrable task representations condense evidence from multiple examples.
-However, they also tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
-  • Models do not appear to condense task knowledge into local task representations in more complicated tasks that require state tracking or chaining multiple subtasks together.
-  • Finally, models can form distinct representations to support generating the same responses when solving tasks independently vs. as part of a broader context.
+These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (davidson2025different; huang2024multimodal; luo2024vision).
+The discovery of transferrable task representations raises several intriguing questions about models’ representations of new tasks.
+How and when are transferrable task representations formed throughout the context?
+How do they differ from other types of task representations that models may use?
+A simple, intuitive hypothesis is that models develop representations for new tasks gradually.
+Task representations might reflect evidence accrual across tokens in the context and refine monotonically into more stable and robust representations.
+This view aligns with the behavioral findings that models perform better with more examples in-context (agarwal2024many; anil2024many).
+We set out to investigate how the dynamics of ICL are reflected in different types of task representations. Our findings suggest a nuanced picture:
+  • We find a stark contrast between two different task representations: an inert representation of task identity is more continuously presented throughout the context, but transferrable task representations only activate sporadically at key tokens.
+  • The fleeting but transferrable task representations can support longer generation. However, their ability to guide model behavior also decays across independent subtask contexts.
+  • Models rely on more distributed task representations in more complicated tasks that require state tracking or chaining multiple subtasks together.
+  • Finally, models form distinct representations of a task when solving it independently vs. as part of a broader context.
 Overall, these results give us a window into language models’ changing state when inferring and solving new tasks in context, but paint a complex and nuanced picture of the dynamics of this state.
@@
 A.
-A schematic of extracting transferrable task representations and restoring task contexts in zero-shot settings.
+A schematic of extracting transferrable task representations and restoring task contexts (via patching) in zero-shot settings.
 The highlighted tokens indicate the source and target for extracting and injecting task representations.
@@
 Transparent bars: task identity decoding accuracy from different token representations.
+2 Related work
+Since the discovery that large language models exhibit emergent in-context learning (brown2020language), there has been substantial interest in investigating this capability and its mechanistic basis.
+From a behavioral perspective, many subsequent works have explored how ICL could develop from implicit meta-learning of data properties (xie2022explanation; chan2022data), and how this may relate to the broader set of language model capabilities (chen2024parallel; lampinen2024broader).
+Some of this work has focused on the surprising fragility of ICL to subtle prompt changes (e.g.
+sclar2024quantifying); conversely, others have highlighted how ICL may be overly robust, allowing “learning” common tasks even if the labels are randomized (min2022rethinking).
+One particularly relevant focus of behavioral work on ICL has been on the dynamics of in-context learning; for example, how adding many example shots can improve performance on difficult tasks, or even those discouraged in post-training (agarwal2024many; anil2024many).
+From a mechanistic perspective, olsson2022context showed in-context learning is supported by induction heads, and other work has studied how they might develop over training (edelman2024evolution; singh2025strategy).
+Extensive ablations have also elucidated core circuits supporting the aggregation of demonstrations from few-shot examples (bakalova2025contextualize; cho2024revisiting).
+Some recent work found that models may create transferrable task representations.
+These are representations that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance.
+hendel2023context demonstrated an instance of such representation: representations at intermediate layers of the last token in few-shot prompts can be injected to mimic the effect of a few-shot prompt.
+Concurrent work from todd2023function identified “function vectors,” which aggregate the effects of multiple attention heads to convey task information.
+Subsequent work has generalized these findings, exploring how function vectors can emerge from instructions (davidson2025different) and how task vectors capture task representations across modalities (huang2024multimodal; luo2024vision).
+Other works have explored how these representations emerge over training (dong2025understanding; yang2025task; yin2025attention), revealed their limitations (dong2025understanding; tikhonov2025one), and extended the methods to more robustly restore task contexts in zero-shot settings (li2024implicit; saglam2025learning).
+Building on these studies, our work characterizes the dynamics of different task representations both within examples in individual tokens and across the context, as well as across different task types.
+We compare and contrast these dynamics and additionally analyze how different task representations co-exist.
 3 Methods
 Tasks
-For our analyses, we built upon tasks from prior work on transferrable task representations (Hendel et al.,, 2023; Todd et al.,, 2024).
+For our analyses, we built upon tasks from prior work on transferrable task representations (hendel2023context; todd2023function).
 The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word or translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words).
@@
 Models
-In the main paper, we present results on the open-weight pre-trained 4B, 12B, and 27B Gemma V3 models (Team et al.,, 2025).
-In Appendix C, we also show that the main findings replicate on the 4B, 8B, and 14B Qwen3 models (Yang et al., 2025a, ).
+In the main paper, we present results on the open-weight pre-trained 4B, 12B, and 27B Gemma V3 models (team2025gemma).
+In Appendix C, we also show that the main findings replicate on the 4B, 8B, and 14B Qwen3 models (qwen3).
+Decoding task identity
+We study a simple representation for new tasks given in-context evidence: whether token representations throughout the context contain robust task identity information.
+We trained simple linear decoders to predict the task category from the layer residual activations of different tokens.
+At each layer and token combination, we used 100 token instances for each task to train and test a task identity decoder.
+All decoders were trained for 20 epochs, with a batch size of 256 and a learning rate of 0.01.
+We report the decoding accuracy across the 25% held-out test representations.
 Extracting transferrable task representations
-We primarily investigate task vectors discovered in Hendel et al., (2023) as a window to study language models’ transferrable task representations.
+We primarily investigate task vectors discovered in hendel2023context as a window to study language models’ transferrable task representations.
+In Appendix D, we show that alternative extraction methods of transferrable task representations like function vectors (see todd2023function) show consistent results.
 We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query, as shown in Figure 1A.
 Task vectors are the layer residual activations extracted from the last token before answer generation (in the example in Figure 1A, this corresponds to the highlighted colon token).
-Hendel et al., (2023) showed that task vectors can reinstate task performance on a different query even without any prior context.
+hendel2023context showed that task vectors can reinstate task performance on a different query even without any prior context.
 Specifically, when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any prior few-shot examples.
-We replicate and extend the procedure outlined in Hendel et al., (2023).
+We replicate and extend the procedure outlined in hendel2023context.
 For each model and task, we first search for the layer that best captures the task representation, using 50 queries from the dataset as the development set and in an 8-shot setting.
 As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations.
-We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model (covering the global attention layers).
+We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; team2025gemma), and every 6 layers starting from layer 5 in the 27B model (covering the global attention layers).
 The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task.
 This best layer is subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset.
-Consistent with prior results, we generally find that task vectors extracted and injected at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
+Consistent with prior results, we generally find that task vectors extracted and patched at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
 Evaluating task transfer
@@
 We show these layer search results across different $k$’s in Figure 2A, but otherwise focus on results from reusing the best layer from the layer search on 8-shot prompts for other $k$’s.
-We then studied whether transferrable task representations form in tokens other than the colon token.
+We then repeat this intervention on other tokens to understand when transferrable task representations form.
 We extracted layer residual token activations for other format tokens in the context, including the “Q”, the “:” following “Q”, the “A”, and the new-line token before “A”.
@@
 All token representations are evaluated on the extent to which they restore task accuracy on zero-shot prompts.
-Decoding task identity
-We also studied whether token representations throughout the context contain robust task identity information.
-We trained simple linear decoders to predict the task category from the layer residual activations of different tokens.
-At each layer and token combination, we used 100 token instances for each task to train and test a task identity decoder.
-All decoders were trained for 20 epochs, with a batch size of 256 and a learning rate of 0.01.
-We report the decoding accuracy across the 25% held-out test representations.
+Lastly, we also explore a set of cross-token analyses to investigate the extent to which different tokens share representational or functional roles (see main text).
+Characterizing the state of task representations.
+We use specific terminology to describe the representational dynamics of the model’s internal state, distinct from the methodological operations (e.g., activation patching) used to probe them.
+By recontextualization, we refer to the phenomena that transferrable task representations reinstate task contexts and enable inference in a zero-shot setting.
+When discussing temporal properties, we refer to the distribution of representations along the sequence dimension.
+Finally, we use task scope to capture the semantic extent of the inference behavior enabled by recontextualization.
 4 Results
-How do language models infer new tasks in-context?
-We leverage the transferability of task contexts to understand how models accrue evidence and refine task representations.
-We find that language models indeed form stronger task representations that aggregate in-context evidence.
-However, this evidence accrual process is non-gradual and happens in a sporadic way.
-In particular, models only form effective task representations that transfer task contexts at certain tokens (Figure 1C, also see Figure 2A).
-This is in strong contrast to a persistent sensitivity of task identity across tokens in the context (Figure 2B).
-We also find that the surprisingly local transferrable task representations tend to only capture a minimal “task scope.” Below, we discuss these findings in more detail.
+How do language models represent new tasks in-context?
+We use the transferability of task contexts to understand when models form transferrable task representations, as opposed to maintaining a simple representation of task identity.
+We first show that transferrable task representations like task vectors indeed aggregate in-context evidence.
+We then show that this evidence accrual process happens in a sporadic way, with transferrable task representations only forming at certain tokens (Figure 1C, also see Figure 2A).
+This is in strong contrast to the persistence of representation for task identity across tokens in the context (Figure 2B).
+We also find that, in different settings, transferrable task representations can support generation for longer tasks or only support a minimal “task scope.” Below, we discuss these findings in more detail.
 Figure 2: Transferrable task representations activate sporadically at key tokens, but task identity representations persist throughout the context.
@@
 See the main text for more details.
-3.1 Local task representations can accrue evidence
+4.1 Transferrable task representations accrue evidence
 Transferrable task representations reflect evidence accrual.
-Consistent with the behavioral gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations also reflect increased task certainty with increased context (Figure 3A).
+Consistent with the behavioral gain from including more examples in-context (e.g., anil2024many), we confirm that transferrable task representations also reflect increased task certainty with increased context (Figure 3A).
 Task vectors extracted following more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples.
@@
 As models appear to successfully accrue evidence in these highly local representations, we sought to understand how the task representations themselves changed over more examples (Figure 4A).
-For this analysis, we look at the task representations extracted from the mode best layer across different tasks.
+We look at the task vectors extracted from the mode best layer across different tasks.
 This is to control for magnitude differences of the residual activations across layers and make a fair comparison.
-Although the best layer for transferrable task representations sometimes differ across tasks, the best layers tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
-In Gemma V3 models, as we increase the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot prompts.
+Although the best layer for transferrable task representations sometimes differ across tasks, the best layers tend to reside in the middle layer range across all model sizes, consistent with prior findings (hendel2023context).
+In Gemma V3 models, as we increase the number of examples, we generally found reduced variance among task vectors extracted from different k-shot prompts.
 This can signal that in-context task representations tend to denoise or converge to more stable representations as models gain evidence.
 The magnitude (L2-norm) of the task vectors also tends to decrease over time.
-However, for both the variance and magnitude, there were considerable differences between tasks and model sizes.
+However, for both the variance and magnitude, there were considerable differences between tasks and models.
 Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2).
@@
 For example, representations for antonym (0), antonym x 3 (14), and where antonym appears as a first task in a mixed-generation task chain (24&27) are close but distinct.
-3.2 Task representations exhibit temporal locality
+4.2 Different task representations exhibit distinct temporal profiles
 The analyses above confirm that transferrable task representations benefit from increasing in-context evidence and converge to better representations.
-To understand the full temporal profile of this accrual process, we repeated the task vector recontextualization experiment on other tokens in the prompt, which revealed that the transferrable task representations do not strengthen monotonically.
+To understand the full temporal profile of this accrual process, we repeated the task vector intervention on other tokens in the prompt, which revealed that these representations do not strengthen monotonically.
+We show that this temporal locality is unique to transferrable task representations, as representations for task identity are widespread.
 Transferrable task representations are not found in most tokens.
-We tested whether extracted representations from other format tokens in the prompt can also restore the corresponding task contexts.
+We tested whether extracted representations from other format tokens can also restore the corresponding task contexts.
 These include “Q”, the “:” following “Q”, “A”, and the new-line token before “A”, which are all shared across examples, tasks, and contexts.
@@
 This suggests that the model is generally task-sensitive, but instantiates transferrable task representations only at particular timepoints in the context.
-3.3 Task representations exhibit scope locality
+Non-trivial cross-token representational and functional transfer.
+We primarily observe task identity representations in all tokens and transferrable task representations in the key pre-answer colon tokens.
+However, we note that these two types of representations are not entirely binary.
+We performed a set of cross-token task vector transfer experiments: both using representations from non-key tokens in few-shot prompts to intervene on the key colon token in zero-shot prompts, and using representations from the key colon token to intervene on non-key tokens.
+More details are presented in Appendix E.
+In both cases, we observe some non-trivial success of task recontextualization.
+For example, representations extracted from the colon following the “Q” token (prior to query presentation) can partially restore task contexts.
+In Qwen3 models, we also find that non-key token sites can functionally participate in task recontextualization when injected with effective transferrable task representations.
+These results suggest that task identity signals and transferrable task knowledge may coexist in token representations.
+Furthermore, token sites that do not usually form transferrable task representations may nonetheless preserve a functional role to establish task contexts.
+We then further explore the overlap between identifiable representations across tokens.
+We show that task identity regressions fit on non-transferrable representations from non-key tokens generalize to transferrable ones, that there is partial overlap between the identifiable-non-transferrable subspaces and the transferrable ones, and that the degree of generalization and overlap is modulated by the same task features as in our other experiments.
+These results reinforce our claim that these different task representations have distinct dynamics and can coexist, but show additional nuance in their partial overlap.
+4.3 Transferrable task representations capture variant scope locality
 We have seen evidence that transferrable task representations tend to be temporally local.
@@
 That is, are the restored task contexts in the zero-shot forward pass also fleeting in nature?
-Prior work has mostly focused on simple, single-token output tasks.
-To study the scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token.
+To study the task scope supported by these representations, we tested to what extent restored contexts can support longer generation beyond the first token.
 Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
@@
 A.
-Bar plot: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks.
-Line plots: recontextualization accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context.
+Bar plot: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks; accuracies within each task are averaged across output units.
+Line plots: recontextualization accuracy for each output unit, conditioned on sequences where models generated full correct responses with eight examples in-context.
 An output unit usually corresponds to a single word and is occasionally a short phrase (e.g.
@@
 Visualization as in A, but for mixed-generation tasks.
-Transferrable task representations tend to support a semantically-independent task scope.
-In these experiments, we find further evidence of the locality of transferrable task representations.
+Transferrable task representations sometimes only support limited task scopes.
+In these experiments, we find further evidence on the semantic locality of transferrable task representations.
 Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 5, bar plots; also see Figure S1).
-Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 5, line plots; see similar results on Qwen3 models in Figure S13), suggesting that the restored task contexts “fade” over longer generation.
-Models seem to be consistently decomposing tasks and forming local task representations that capture a minimal “task scope”.
-For example, in choose_first_middle_last_of_5, the reinstantiated task context only supported generating the first word, suggesting that further token generations rely on additional representations instantiated by a separate mechanism while processing the previously-generated tokens.
-This effect is very pronounced in the mixed-generation tasks.
-In these cases, all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain, and defer representing later subtasks.
-Interestingly, the extracted task representations for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 4B), even though they support generating the same responses.
-This can potentially reflect an inert sensitivity to the broader mixed-task context based on vocabulary differences.
-Additionally, these differences may also convey information relevant for later subtasks to be activated, even though they are not sufficient to directly restore execution of later subtasks.
-The restored task contexts do support longer generation to some extent in many tasks, including repeating a simple task multiple times and some list-level tasks (i.e., reversing or shifting a word list).
-Overall, these results suggest that language models may be automatically segmenting semantically-independent task scopes when possible, such that task representations for longer or composite tasks are offloaded onto multiple tokens.
-4 Related work
-Since the discovery that large language models exhibit emergent in-context learning (Brown et al.,, 2020), there has been substantial interest in investigating this capability and its mechanistic basis.
-From a behavioral perspective, many subsequent works have explored how ICL could develop from implicit meta-learning of data properties (Xie et al.,, 2022; Chan et al.,, 2022), and how this may relate to the broader set of language model capabilities (Chen et al.,, 2024; Lampinen et al.,, 2024).
-Some of this work has focused on the surprising fragility of ICL to subtle prompt changes (e.g.
-Sclar et al.,, 2024); conversely, others have highlighted how ICL may be overly robust, allowing “learning” common tasks even if the labels are randomized (Min et al.,, 2022).
-One particularly relevant focus of behavioral work on ICL has been on the dynamics of in-context learning; for example, how adding many example shots can improve performance on difficult tasks, or even those discouraged in post-training (Agarwal et al.,, 2024; Anil et al.,, 2024).
-From a mechanistic perspective, Olsson et al., (2022) showed how induction heads could support in-context learning, and other work has studied how they might develop over training (Edelman et al.,, 2024; Singh et al.,, 2025).
-More recently, attention has turned to cases where models may create internal task representations.
-These are representations that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance.
-Hendel et al., (2023) demonstrated an instance of such task representations: representations at intermediate layers of the model at a key token that can be injected to mimic the effect of a few-shot prompt.
-Concurrent work from Todd et al., (2024) identified “function vectors,” which aggregate the effects of multiple attention heads that convey task information.
-Subsequent work has generalized and extended these findings, for example exploring how function vectors can emerge from instructions as well as examples (Davidson et al.,, 2025) and how task vectors capture task representations in multimodal models (Huang et al.,, 2024; Luo et al.,, 2024).
-Other works have explored how these representations and structures emerge over training (Yang et al., 2025b, ; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024).
-Our work builds on these findings and use transferrable task contexts as a window to study the dynamics of in-context task representations in language models.
+Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases over longer generation (Figure 5, line plots; see similar results on Qwen3 models in Figure S13), suggesting that the restored task contexts sometimes “fade” over longer generation.
+We notice a strong contrast between different tasks.
+In some longer-generation tasks that repeat the same task multiple times or require list-level operations, the reinstantiated task contexts (from task vectors extracted from and patched onto the key colon token) successfully supported continued generation.
+In other tasks, however, this intervention only supported generating the first output.
+This effect is very pronounced in the mixed-generation tasks that combine multiple distinct subtasks.
+In these cases, all models form strong local task representations that only encapsulate the first subtask.
+This suggests that transferrable task representations at the key colon token can capture variant “task scopes”, which tend to be limited in cases where the subtasks have distinct semantic boundaries.
+We confirm that models defer representing later subtasks using additional interventions (see Appendix F).
+Indeed, we find that transferrable subtask representations in mixed-generation tasks form at the comma tokens prior to subtask answer generation.
+However, intervening comma tokens alone does not seem to recover the full subtask inference ability.
+We posit that the full subtask representation is distributed across colon and comma tokens.
+Interestingly, the extracted task representations from the colon tokens in the same simple task can be distinct when it appears independently or as a first task in a multi-task context (Figure 4B), even though they support generating the same responses.
+This may in-part reflect vocabulary differences in later subtasks, but potentially also information relevant for later subtasks to be activated, even though these representations alone are not sufficient to restore execution of later subtasks.
 5 Discussion
 We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks.
-Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task contexts in zero-shot settings.
-Our results show that in many tasks, models refine task representations over more evidence such that the representations more successfully restore task contexts.
-However, these transferrable task representations only sporadically activate, and seem to best support a minimal task scope (e.g., a first irreducible subtask).
-The dynamics of these effective, transferrable task representations strongly contrast with a general sensitivity to high-level task differences that persists throughout the context.
-In addition, we find many cases where models do not appear to condense global task representations into a local, token-level representation, such as tasks involving more state tracking and tasks chaining different types of subtasks together.
-In general, our results complicate the intuitive picture that the computational process underlying language models’ in-context learning and task inference is smooth and gradual.
-Models do not appear to refine task representations at a per-token basis, not even in a step-wise manner—even on a per-example basis, the task state is not sustained, but often fades and reactivates across different tokens.
-That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information may suggest a general inductive bias to compress knowledge to local representations when possible.
-This may also relate to the success many works have observed on extracting transferrable task representations in broad settings, such as following instructions or images (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024), or capturing information for multiple possible task outcomes (Xiong et al.,, 2024).
-However, one trend that arose from these investigations is that language models do not seem able to condense task information into local representations in all cases.
-Rather, they exhibit a tendency to form sharp local contexts for small task units, and offload broader task contexts such as mixed or multiple tasks across time.
-As discussed earlier, even for simple tasks where some intermediate-state tracking may be required, successful inference may need to rely on cross-token and cross-layer computation (e.g., as shown in Ameisen et al.,, 2025).
-In these cases, the effective restoration of the computation process may also require intervening multiple components during the forward pass.
-It is also possible that by overriding token activations at intermediate layers with task vectors, the models have lost any task state information formed in earlier layers.
-Other methods that restore task states through additive injection rather than activation patching may therefore be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024).
-We also note that different models may differ in the extent to which they form local or distributed effective representations across the context or layers.
-An intriguing direction for future work is to study the mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations.
+We evaluated when in the context we can extract transferrable task representations that restore task contexts in zero-shot settings, and when we can extract robust representations of task identity.
+Our results show that, transferrable task representations only sporadically activate, even though they accrue evidence from multiple examples.
+However, models maintain a strong general sensitivity to high-level task differences that persists throughout the context.
+We find cases where models do not form a global task representation at particular token sites, such as tasks involving state tracking and tasks combining distinct subtasks together.
+We also find nuances in the distinction between the two types of task representations, where the representational and functional roles of different tokens exhibit non-trivial overlap in certain settings.
+In general, our results complicate the intuitive picture that language models smoothly and gradually refine task representations during in-context learning (ICL).
+Models form task representations at different levels and different rates.
+Across tokens and examples in the context, the same task state is not sustained, but can fade and reactivate.
+These high-level changes in models’ representational states unite a few observations from prior work (bakalova2025contextualize; cho2024revisiting; dong2025understanding).
+Key tokens prior to answer generation help aggregate in-context examples but do not all participate in prediction circuits, as the effective task representations they once generate are deactivated and reactivated later.
+The convergence of transferrable task representations may result from the contextualization subcircuit identified in (bakalova2025contextualize), such that later demonstrations contribute more to the test query as seen in (cho2024revisiting).
+Tokens seemingly outside of a core ICL circuit may nonetheless help restore task contexts to some extent, potentially contributing to the bypass mechanisms noted in (cho2024revisiting).
+One trend that arose from these investigations is that language models do not seem to condense task information into local representations in all cases.
+We find that language models can construct token-level task representations that flexibly capture more broad task contexts and longer outputs.
+But in some cases, models elect to form sharp local contexts only for small task scopes and offload broader task representations across multiple tokens.
+This flexibility across different types of longer-generation tasks adds additional nuance to the distributed-ness of effective task vectors across multiple tokens similarly noted in dong2025understanding and tikhonov2025one.
+This flexibility may also relate to the success many studies have observed on extracting transferrable task representations in broad settings, such as following instructions or images (davidson2025different; huang2024multimodal; luo2024vision), or capturing information for multiple possible task outcomes (xiong2024everything).
+For some tasks that require more complicated computation such intermediate-state tracking, successful inference may need to rely on not only cross-token but cross-layer representations (e.g., as shown in ameisen2025circuit).
+In these cases, the effective restoration of the computation process may also require intervening multiple layers during the forward pass.
+It is also possible that by overriding token activations at intermediate layers with task vectors, the models have lost task state information formed in earlier layers.
+Here, we show that zero-shot recontextualization remains challenging in these tasks even through additive injection such as function vectors (todd2023function, see Appendix D).
+But more advanced methods that restore task states by intervening multiple layers may be more successful at restoring model task states in these cases and elicit different temporal or semantic dynamics (li2024implicit; saglam2025learning).
+We also show some evidence that different models may differ in the extent to which they form local or distributed representations in different tasks.
+An intriguing direction for future work is to study the mechanistic bases for the strong temporal and scope locality we observed in models’ in-context transferrable task representations.
 One possibility may be that the residual stream is more stable and easier to learn from during training.
@@
 Some of these features may even be exclusive to models pre-trained on natural languages (e.g.
-Yang et al., 2025b, ).
+yang2025task).
+Taken together, our findings offer some useful insights for interpretability and interactions with language models at large.
+For the science of understanding language models, comparing between more passive and more active, transferrable levels of representation offers a useful framework for probing representations of different knowledge or concepts.
+Practically, the variant temporal and semantic locality of different task representations means that interactions such as prompting or model steering should not assume a stable, continuous task state.
+Rather, effective control may require strategic re-instantiation of the fleeting representations at critical subtask boundaries.
 Limitation
 We note a few important limitations of our work.
-First, we primarily rely on task vectors as the method to extract transferrable task representations.
-This means that our conclusions and speculations are bounded by the effectiveness of this method.
+First, we primarily rely on single-token, single-layer intervention methods to extract transferrable task representations.
+This means that our conclusions and speculations are bounded by the effectiveness of these methods.
 As we discussed earlier, representations for some task contexts may be more distributed, either across tokens and/or across model layers.
-It would be important to confirm if similar dynamics are observed in less-constrained methods such as a multi-layer recontextualization (Li et al.,, 2024).
+It would be important to confirm if similar dynamics are observed in less-constrained methods such as a multi-layer recontextualization (li2024implicit; saglam2025learning).
 Second, we mostly explored relatively simple tasks, including when we investigated longer-generation tasks.
-It’s possible that many of the dynamics we observe here would not generalize to settings with naturalistic languages, especially when the tasks are not so cleanly decomposable and a single, semantically-independent task unit is hard to define.
+It’s possible that many of the dynamics we observe here would not generalize to settings with naturalistic languages, especially when the tasks are not cleanly decomposable and a single, semantically-independent task scope is hard to define.
 Conclusion
 We investigated how the dynamics of in-context learning are reflected in the development of language models’ internal task representations.
-Our results suggest that language models do not smoothly refine a global task state in-context.
-While general task sensitivity persists throughout context, models appear to construct effective task representations in a “just-in-time” fashion to solve a simple task scope immediately ahead.
-The fleeting, minimally-scoped nature of these in-context task representations provides new insight into the models’ state of inferring and performing tasks based on new evidence.
+Our results suggest that language models do not smoothly refine a global task state.
+While general task sensitivity persists throughout the context, models construct more active, transferrable task representations in a “just-in-time” fashion.
+These contrasting levels of task representations provide new insights into the models’ state of inferring and performing tasks based on new evidence.
 Acknowledgments and Disclosure of Funding
-We thank Yasaman Bahri, Katherine Hermann, Mike Mozer, Sjoerd van Steenkiste, and Murray Shanahan for helpful comments and support on this work.
+We thank Yasaman Bahri, Katherine Hermann, Mike Mozer, Sjoerd van Steenkiste, Murray Shanahan, and anonymous reviewers for helpful comments on this work.
 Appendix A Tasks
-Table 1: Simple/shorter-answer tasks. See Todd et al., (2024) for more details for the first nine tasks.
+Table 1: Simple/shorter-answer tasks. See todd2023function for more details for the first nine tasks.
 Task Name Example
@@
 Figure S13: In Qwen3, restored task contexts in longer- and mixed-generation tasks also decay over more output units during generation. Visualization as in Figure 5.
+Appendix D Temporal and semantic locality in Function Vectors
+We show that the main locality features of transferrable task representations replicate when using function vectors (todd2023function) as the extraction and recontextualization method.
+Function vectors are the summed activations of a set of critical attention heads that contribute to in-context task performance.
+Similar to task vectors, function vectors can transport task semantics across a range of tasks when additively injected onto the last token in a zero-shot prompt.
+We build on the procedure outlined in todd2023function to extract function vectors from the key last “:” token in 8-shot prompts, with the following exception: we separately measure the critical attention heads contributing to a particular task and extract a per-task function vector.
+We use the top 16 heads for QWEN3_4B and GEMMA_V3_4B, 32 heads for QWEN3_8B, 48 heads for GEMMA_V3_12B, 64 heads for QWEN3_14B, and 80 heads for GEMMA_V3_27B, roughly corresponding to the number of critical attention heads used in similar-sized models in todd2023function.
+To remain comparable with the task vector extraction procedure mentioned above, we similarly sampled at most 50 queries per task to extract function vectors and search for the best intervention layer in zero-shot prompts.
+We then test the function vector with the best-layer intervention on the remaining queries.
+In the function vector case, the development set queries are sampled from cases where models correctly generate the target answer in an 8-shot prompt.
+Finally, we generalize the extraction of function vectors on other tokens in the prompt to study when this kind of transferrable task representation form.
+We re-compute the critical attention heads and the best intervention layer per token site.
+Figures S15 and  S16 show replication of the main temporal and semantic scope locality of transferrable function vector representations on Qwen3 models.
+These results suggest that function vectors and task vectors extract similar kinds of transferrable task representations during models’ in-context learning.
+We unfortunately have yet to be able to induce zero-shot task recontextualization using function vectors on Gemma3 models.
+We suspect that Gemma3 models’ normalization scheme and the increasing magnitude of their residual streams may require scaling up the signals from function vectors when intervening on later layers, but this is a space requiring more investigation.
+Figure S14: Recontextualized zero-shot task performance using function vectors.
+Figure S15: Effective function vectors, like task vectors, tend to only activate on the key colon tokens.
+The plot shows the zero-shot accuracies when the models are recontextualized with the per-task function vector. Accuracies are averaged across the 14 simple tasks.
+Figure S16: Function vectors extracted from the key colon token similarly capture limited scope locality, especially in cases where the overall task combines semantically independent subtasks.
+Visualization as in Figure 5.
+Appendix E Cross-token analyses
+We conduct a set of cross-token transfer and identification experiments to further understand the representational and functional roles of non-key tokens, and the overlap between identifiable and transferable task representations.
+Cross-token task vector transfer We first test two variants of cross-token task vector intervention.
+We use representations from non-key tokens in few-shot prompts to intervene the key “:” token in zero-shot prompts (Figure S17).
+We also test using representations from the key “:” token in few-shot prompts to intervene different token sites in zero-shot prompts (Figure S18).
+For each token pair in either experiment, we re-compute the most effective layer of the cross-token intervention.
+In both cases, we observe some non-trivial success of task recontextualization.
+Although there is notable task and model dependence, we find that representations of the “:” token after the “Q” token (prior to query presentation) can be used to help restore task contexts in zero-shot settings.
+We also find that in Qwen3 models, multiple non-key token sites can functionally participate in task recontextualization when given effective transferrable task representations.
+These results suggest a degree of shared task identity representation and transferrable task representation in some tokens.
+They also indicate that token sites outside of a core circuit for in-context learning (ICL) may nonetheless preserve a useful functional role when given the right information.
+This could potentially contribute to bypass mechanisms noted in (cho2024revisiting), where some residual ICL ability remains in the model even when ablating core ICL circuits.
+Figure S17: Recontextualized zero-shot accuracy across tasks, using representations from different tokens in 8-shot prompts to intervene the last “:” token in zero-shot prompts.
+The x-axis indicates the token representations used to intervene zero-shot inference.
+Each row represents one task.
+We exclude samples where the correct output is simply the first word of the input query.
+Figure S18: Recontextualized zero-shot accuracy across tasks, using representations from the last “:” token in 8-shot prompts to intervene different tokens in zero-shot prompts.
+The x-axis indicates the token site intervened during zero-shot inference.
+Each row represents one task.
+We exclude samples where the correct output is simply the first word of the input query.
+Analyzing how task identity representations overlap with transferable ones at key tokens: In this section, we further analyze the relationship between the identifiable and transferable task representations.
+Specifically, we fit linear classifiers that extract identifiable task representations across all the non-effective format tokens—i.e., all format tokens except the 2nd ‘:’ where we find transferable task representations.
+We also test a version without the first ‘:’ token, given the results above.
+We then test how well these classifiers generalize to identifying the task representations on the held-out 2nd ‘:’ token representations from a disjoint set of prompts.
+We find that there is reliable generalization (Table 4; Figure S19).
+In fact, we observe perfect generalization for all tasks involving a single output, as well as some of the mixed-generation tasks such as antonym + product-company.
+However, for many list-operations tasks (e.g.
+reverse_all_of_3), results are somewhat less reliable; but the classifier still achieves robust 75-80% generalization (far above chance performance of 3%, or 6% among these tasks).
+The performance is comparable with or without the first ‘:’ in the training set for the classifier, suggesting that this generalization is not primarily due to the partially-transferable representations on that token observed above.
+Together, these results suggest that identifiable task representations persist in a relatively consistent way across both tokens that do not produce transferable task representations, and those that do produce transferable task representations.
+This suggest that perhaps the identifiable and transferable task representations lie in subspaces that are not fully overlapping.
+To understand these results more deeply, we analyzed the relationship between the linear dimensions identified by the regression, and the dimensions of variation in the transferable task representations (Figure S20).
+We performed PCA over the transferable representations, and identified how strongly the regression features projected onto the top 1-20 principle components.
+Consistent with the results above, we found that for the same cluster of simple and non-list-operations tasks for which the regressions generalized best, the regression also projected fairly strongly onto the top principle components (with 40%-60% of the identifiable dimension projecting onto the top 20 principle components of transferable representations).
+By contrast, identifying the other list-operation tasks generally seemed to rely more heavily on components outside the top PCs, with only 10-25% of the identifiable dimension projecting onto the top 20 principle components of the transferable representations.
+Correspondingly, we find that if we project the transferable representations down to the top 20 principle components, we find strong generalization (99.1%) for the simpler tasks compared to weaker generalization for the other list-operations tasks (34.8%).
+Together, these results suggest that identifiable and transferable task representations lie in partially-overlapping subspaces of the models representations, but the degree of that overlap is modulated by the task type in similar ways to our other findings.
+These results reinforce our suggestion that these distinct types of task representation are distributed differently across the sequence, but shed more light on their interactions and their overlap.
+Table 4: Task-identifying regressions generalize from non-transferable tokens to transferable ones.
+Training tokens Test tokens Task set Generalization Top-20-PCs gen.
+[…]
+Other tasks 1.0 1.0
+[…]
+Other tasks 1.0 -
+Figure S19: The confusion matrix for the task-identifying regressions (when tested on generalization from non-transferable to transferable representations) shows more details of the errors—while many tasks are identified perfectly, some of the list operation tasks yield more errors.
+Nevertheless, generalization performance is relatively high even among the list-operations tasks.
+Figure S20: Identifiable task representations partly overlap with the top principle components of transferable task representations, but the degree of this overlap is modulated by the task type.
+Additional details: We performed these experiments in layer 29 of Gemma_V3_27B_PT, which was identified as the mode best layer for transferable task representations.
+We used 8-shot prompts, and fit the linear classifiers via $\ell^{2}$-regularized logistic regression.
+We also fit regressions within only the shared-vocab tasks used in some of the analyses above, and we observe qualitatively similar patterns of results in that setting.
+Appendix F Inspecting subtask representations in mixed-generation tasks
+We showed above that, in mixed-generation tasks, recontextualizing zero-shot inference with task vectors extracted from the key “:” was not enough to support continued generation of subsequent subtask answers.
+Here, we explore two hypotheses of how subtask representations may be restored through intervention.
+We take representations from the “:” token in 8-shot prompts and intervene at the “,” tokens prior to subtask answers in zero-shot prompts.
+We compare this intervention to using representations from the matching “,” tokens prior to subtask answers in 8-shot prompts to intervene the “,” tokens during zero-shot answer generation.
+Table 5 demonstrates the intervention success when “,” tokens prior to subtask generation are intervened with representations from matching prior “,” tokens, rather than representations from “:” tokens.
+The results are consistent with findings from tikhonov2025one, suggesting that the formation of transferrable subtask representations can be delayed to later format tokens, especially in cases where the subtasks are semantically independent.
+However, we also note that the restored subtask accuracies are still below that of single-task recontextualization.
+This may reflect that the full subtask representation requires a combination of task representation in “:” tokens as well as in “,” tokens.
+Table 5: Interventions on comma tokens in mixed-generation tasks.
+The table shows the intervened zero-shot accuracy for the output immediately following the intervened “,” token versus any remaining output.
+Results are averaged across tasks.
+Source token Model Output after intervened “,” Remaining Output
+“:” Gemma_V3_4B_PT 0.004 0.000
+Gemma_V3_12B_PT 0.004 0.003
+Gemma_V3_27B_PT 0.003 0.014
+Qwen3_4B 0.003 0.000
+Qwen3_8B 0.001 0.000
+Qwen3_14B 0.004 0.009
+matched “,” Gemma_V3_4B_PT 0.455 0.002
+Gemma_V3_12B_PT 0.295 0.001
+Gemma_V3_27B_PT 0.413 0.007
+Qwen3_4B 0.198 0.003
+Qwen3_8B 0.293 0.011
+Qwen3_14B 0.248 0.016